OpenStudy (anonymous):

Given the two expressions shown below: square root of 4 plus square root of 25 square root of 4 plus square root of 9 Which statement best describes the two expressions? Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational.

OpenStudy (anonymous):

please help ill attach the image so you can understand better

OpenStudy (anonymous):

both are rational this is because the square root of 4 is a real valued integer and so is the square root of 25, and 9 if you add 2 + 5 you get 7 if you add 2 + 3 you get 5 the purpose of this exercise is to test whether or not you know what happens when you take the square root of a number example: the square root of 3 is an irrational number if you add sqrt(3) to 1 the answer is irrational if you add sqrt(3) to ANY integer(1,2,3,4,5...) the answer is irrational

OpenStudy (anonymous):

thanks alot do you think you could help with another question

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

Simplify: cube root of 5 over fourth root of 5 5 to the power of 1 over 4 5 to the power of 1 over 12 5 to the power of 7 over 12 5 to the power of 4 over 3

OpenStudy (anonymous):

ill show you the question

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

your first step is going to be to rationalize the denominator so you will multiply both the top and the bottom by 5^(3/4) the reason for this is because in your denominator you have 5^(1/4), which is irrational 5^(3/4) * 5^(1/4) = 5 this occurs because when multiplying numbers with exponents, you add the exponents the expression now contains your numerator 5^(1/3) * 5^(3/4) and denominator of 5 where the exponents will simplify by the following: 1/3 + 3/4 - 1 = 4/12 + 9/12 - 12/12 = 13/12 - 12/12 = 1/12 so 5^(1/12) is your answer, or the twelfth root of 5

OpenStudy (anonymous):

thanks um do you know how to work with function cuz i need help on this one question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok here are the questions thanks for your help

OpenStudy (anonymous):

system 2 and 4 because when you add 4x - 5y = 2 2 * (3x - 2y = 1) you obtain 10x - 9y = 4 which is the second equation in system 4

OpenStudy (anonymous):

thank you that makes sense

OpenStudy (anonymous):

for the next question to find the y intercept of both f(x) and g(x) you evaluate the function when x is zero this is quite simple: f(0) = -6(1.02)^(0) f(0) = -6(1) ( because any number other than 0 to the power of 0 is equal to the integer 1) f(0) = -6 and from the table you can obtain g(0) g(0) = -3 it seems the y intercept of f(x) is 2 times the y intercept of g(x)

OpenStudy (anonymous):

thankz so much do you think you could help with this last question

OpenStudy (anonymous):

sure no problem

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

your value of g(x) is constant, so it would not depend on x at all so g(x) = 4 your value of f(x) is increasing at an exponential rate and seems to depend on x f(1) = 3 and 3^1 = 3 f(2) = 9 and 3^2 = 9 f(3) = 27 and 3^3 = 27 f(4) = 81 and 3^4 = 81 f(5) = 243 and 3^5 = 243 it seems that defining this function as 3^x = f(x) would work nicely so f(x) + g(x) = 3^x + 4

OpenStudy (anonymous):

thanks you so much for all your help your a lifesaver merry christmas

OpenStudy (anonymous):

:) merry christmas